OpenStudy (anonymous):

21/25 * 10/99 *27/14 As a fraction in lowest terms???

OpenStudy (tkhunny):

Ever hear of a Prime Factorization?

OpenStudy (anonymous):

how do I do that?

OpenStudy (tkhunny):

Take each value and factor it until you have onlt prime numbers remaining. I'll do one. You do the other five. 99 = 9*11 = 3*3*11 Go!

OpenStudy (anonymous):

I'm lost

OpenStudy (tkhunny):

Okay, I'll do another one. You do the other four. 99 = 9*11 = 3*3*11 25 = 5*5 21 = 10 = 27 = 14 = Fill it in!

OpenStudy (tkhunny):

Why did you close the question? You have prime factorizations to do!

OpenStudy (anonymous):

Working on them LOL